# Comparison of generalised additive models and neural networks in applications: A systematic review

### v1

@@ 6.1 Paper-level characteristics
Table 4: Papers included in the systematic review.
Domain area Papers Datasets
Environmental & Agricultural Sciences 66 199
Aguilos et al., (2024), Beaugrand, (2024), Bonsoms and Ninyerola, (2024), Fan et al., (2024), Fu et al., (2024), Gazis et al., (2024), Jung et al., (2024), Lei et al., (2024), Mouta et al., (2024), Senaratna et al., (2024), Yang et al., (2024), Ghorbani et al., (2024), Adamova and Ukrainskiy, (2023), Boudreault et al., (2023), DelSanto et al., (2023), Li et al., (2023), Yan et al., (2023), Benemann et al., (2022), Shafizadeh-Moghadam et al., (2022), Yadav et al., (2022), Evans and Shen, (2021), Ma et al., (2021), Martín et al., (2021), Mondal and Bhat, (2021), Phelps and Woolford, (2021), Tian et al., (2021), Zhang et al., (2021), Dodangeh et al., (2020), Dorich et al., (2020), Mandal et al., (2020), Yadav et al., (2020), Mugo and Saitoh, (2020), Nguyen et al., (2019), Liu et al., (2018), Pourghasemi and Rahmati, (2018), Corona-Núñez et al., (2017), Gao et al., (2017), Pôças et al., (2017), Ouarda et al., (2016), Shortridge et al., (2016), Vogel and Brown, (2016), Durocher et al., (2015), Li et al., (2015), Lin et al., (2015), Xu et al., (2014), Gutiérrez-Estrada et al., (2012), Queirolo et al., (2012), Haywood and Stone, (2011), Palialexis et al., (2011), Aertsen et al., (2010), Gutiérrez-Estrada and Bilton, (2010), Ogawa-Onishi et al., (2010), Tisseuil et al., (2010), Virkkala et al., (2010), Gutiérrez-Estrada et al., (2009), Lee et al., (2009), López-Darias et al., (2008), Araújo et al., (2005), Luoto and Hjort, (2005), Megrey et al., (2005), Wang et al., (2005), Thuiller, (2003), Brosse and Lek, (2002), Brosse and Lek, (2000), Brosse et al., (1999), Shatar and Mcbratney, (1999)
Engineering, Mathematics, & Computer Science 35 141
Chen et al., (2024), Kang and Choi, (2024), Nascimento et al., (2024), Olca and Yücel, (2024), Pachauri, (2024), Shojaeian et al., (2024), Berrisch et al., (2023), Koh et al., (2023), Ledmaoui et al., (2023), Montero et al., (2023), Razaq and Shukur, (2023), Zhang and Yang, (2023), Zhang et al., (2023), Duan et al., (2022), Mestanza et al., (2022), Pachauri and Ahn, (2022), Dudek, (2021), Fang et al., (2021), Kamis et al., (2021), Matsumoto and Yamada, (2021), Yang et al., (2021), Daniel et al., (2020), Park et al., (2020), Ilseven and Gol, (2019), Tsang et al., (2018), Ilseven and Gol, (2017), Martín et al., (2016), Francisco-Fernández et al., (2015), Yang et al., (2015), Fukuda et al., (2013), Aertsen et al., (2011), Berg, (2007), Schlink et al., (2006), Morlini and Orlandini, (2001), Faraway and Chatfield, (1998)
Health & Biomedical Sciences 12 25
Francisco et al., (2024), Esmaeili et al., (2023), Xu et al., (2023), Moslehi et al., (2022), Shoko et al., (2022), Taylor et al., (2019), Allard et al., (2017), Cengiz and Terzi, (2012), Gregori et al., (2011), Gregori et al., (2009), Giachino et al., (2007), Goldfarb-Rumyantzev and Pappas, (2002)
Other 30 65
Pasha and Reddy, (2024), Roldán et al., (2024), Szeląg et al., (2024), Amara-Ouali et al., (2023), Balan et al., (2023), Bolleddula et al., (2023), Cai et al., (2023), Krämer et al., (2023), Mathur and Mathur, (2023), Sanii and Chan, (2022), Zhou and Zhang, (2022), Dwivedi et al., (2021), Fotiadis et al., (2021), Frasier et al., (2021), Li et al., (2021), Rengarajan et al., (2021), Zhu et al., (2021), Anglart et al., (2020), Zhu et al., (2020), Zihan et al., (2020), Wheeler et al., (2019), Baquero et al., (2018), Luan et al., (2018), Davoudi et al., (2017), Nyhan et al., (2014), Liu and Tesfamariam, (2012), Ryser and Denzler, (2009), Marmion et al., (2008), Heikkinen et al., (2007), Lee et al., (2002)
@@ 6.2 Dataset-level characteristics
Feedforward Multilayer perceptron (MLP) neural networks were the most common type, accounting for 89% of datasets. All other neural networks were variations of this, and any one type accounted for less than $5\%$ of the datasets. Neural network complexity was calculated as the total number of parameters, excluding those in the input layer. This exclusion was necessary because input nodes correspond to the number of predictors, which is already captured by the fact that we have collected the number of predictor variables as an attribute. As mentioned, neural network complexity is among the data items frequently unreported. Only 35% of datasets provided sufficient detail to compute the full parameter count. However, 64% reported the number of hidden layers, which includes instances where layers were discussed but not the number of nodes in each needed to look at full complexity. Depth of a neural network via the number of hidden layers is, in its own right, a feature of interest. Among the datasets that did report the number of hidden layers, 70% used a single hidden layer, 22% had two hidden layers, and 7% used between three and five. This distribution indicates that the majority of neural networks used in comparisons with GAMs were shallow architectures, typically with one or two hidden layers.
@@
Looking at author-reported outcomes, neural networks were declared better than GAMs in 27% of datasets. This is more than the 19% of datasets in which GAMs were reported as better than neural networks. In 9% of datasets, both model types were found equivalent. However, author-reported outcomes provide an incomplete and often inconsistent comparison between GAMs and neural networks. Differences can arise in how outcomes are classified from author to author. Additionally, 46% of author outcomes are undefined and hence recorded as “unclear”, often due to a focus on an alternative model in that application. A way to address these limitations is to directly analyse the performance metric values reported for both GAMs and neural networks in each instance. This allows for a more objective comparison and quantifies the magnitude of performance differences, which is not possible when looking at the qualitative author outcomes alone.
@@ 7 Performance metric analysis
To quantitatively synthesize the comparative performance of GAMs and neural networks across the reviewed literature, we create a statistical model for the reported performance metric values directly. This analysis examines how the relative performance, measured by the most frequently used metrics, is associated with key study characteristics such as sample size, domain area, and model complexity. Given the hierarchical structure of the data we have collected, where datasets are nested within papers, a mixed-effects modelling approach was employed to account for intra-paper dependence.
@@ 7.1 Methods
Linear mixed-effects models were fitted using the lme4 package (Bates et al.,, 2015) in R programming language (R Core Team,, 2024). We included a paper-level random intercept effect to account for correlated results within a given paper, along with the following fixed effects: publication year, domain area, sample size, number of predictors, and neural network complexity. Pairwise interactions with year of publication were also included to account for trends over time.
7.1.1 Response and predictor variables
As shown in Table 8, RMSE, $R^{2}$, and AUC are, by far, the most commonly-used performance metrics, accounting for 53%, 33%, and 26% of datasets respectively and collectively covering 89% of all datasets. RMSE and $R^{2}$ are common performance metrics used in regression, accounting for 76% and 48% of regression datasets respectively, and 91% collectively. AUC, a standard metric for classification models, was used in 82% of our classification datasets. These three metrics were selected for our performance metric analysis due to their prevalence and large coverage of datasets.
@@ 7.1.1 Response and predictor variables
where positive values indicate superior neural network performance for that metric, negative values signify GAM performance was better, and zero indicates model equivalence. For AUC and $R^{2}$, they are first converted to losses, in which 0 indicates optimal performance, by subtracting them from 1. This allows for all three to represent loss ratios. The ratio of performance metrics provides a relative performance measure that is invariant to scale. This is necessary here, as we are comparing results across many different datasets with substantial variation in absolute performance; different datasets lie on different scales, with differing levels of achievable performance given the inherent noise/uncertainty in the problem at hand. Moreover, we have taken the logarithm to place the response on the real line (so that it is suitable for linear modelling) and to address some skewness. Figure 5 displays the log performance ratios for datasets using RMSE, AUC, and $R^{2}$; points are coloured based on the performance metric used. The horizontal line at zero indicates where GAMs and neural networks are equivalent; points above the line correspond to datasets where neural networks outperform GAMs, while points below are where GAMs achieve superior performance. Overall, we can see that there is a high level of variability in results, with no firm evidence in favour of neural networks or additive models (as the points are generally centred around zero), and no strong trends over time. However, this high-level overview of results does not account for predictor variables (other than the year) or the random effects of papers.
@@
The following predictor variables were considered in the meta-analysis as fixed effects: year of publication (Year), domain area (Area), sample size (Sample_Size), number of predictors (No_Predictors), and neural network complexity (Complexity). Paper ID (Paper_ID) is used as a random effect to account for the dependence between datasets coming from the same papers. To retain as much information as possible, numerical predictors with missing values were converted to categorical variables. Tertiles of the numerical variables were created along with a category “Unknown”, resulting in four levels. This approach enabled the inclusion of incomplete records and allowed assessment of whether missingness was associated with performance outcomes.
Year (numerical, complete):
Values were centred by subtracting the middle year, 2011, from each observation (range: 1998–2024).
Area (categorical, complete):
Classification based on the Scopus categories further grouped into four overarching areas.
Sample_Size (numerical, incomplete):
@@
The number of observations in each dataset. Categorical levels: Small $\leq 350$, Medium (350 - 2,800), Large $\geq 2,800$, and Unknown.
No_Predictors (numerical, incomplete):
@@
The number of predictors in each dataset. Categorical levels: Small $\leq 5$, Medium (5 - 10), Large $\geq 10$, and Unknown.
Complexity (numerical, incomplete):
@@
Number of parameters in the neural network, excluding the input layer. Categorical levels: Small $\leq 6$, Medium (6 - 150), Large $\geq 150$, and Unknown.
Interactions of the predictor variables with year were explored to assess whether or not the effects of these variables change through time. For all models, the category “small” is used as the reference level and 2011 (transformed to be zero) as the reference year.
7.1.2 RMSE metric analysis
@@
The RMSE analysis included 229 dataset observations from 80 papers. After removing two outliers, 227 dataset observations, from 78 papers remained. Including covariates in the analysis allows us to explain more of the variation in RMSE ratios across datasets, and to investigate how aspects of the datasets and models influence the relative performance of neural networks and GAMs. The RMSE model estimates and standard errors (SE) are presented in Table 9. (Models for AUC and $R^{2}$ are also shown, but discussed in Section 7.1.3 and Section 7.1.4 respectively.) The model included all fixed-effects predictors and their pairwise interactions with time (Year) as follows:
@@
where $(1|\texttt{Paper\_ID})$ represents a random intercept capturing variation among datasets. Note that modelling the ratio in this manner is a special case of a slightly more general model discussed in Appendix A.
The intercept for this model is positive, however, not statistically significantly different from zero. This indicates that, at the reference levels for covariates, there is no evidence of a distinction between the performance of GAMs and neural networks. Neural network complexity and its interaction with time significantly affect the log RMSE ratio. From the model output, medium and large complexity effects are both negative and significant, indicating that, in 2011, more complex neural networks performed worse than GAMs compared to smaller neural networks. The complexity-year interactions for medium and large neural networks are positive and significant; this indicates that larger neural networks have been improving over time relative to small neural networks.
Figure 6 displays the marginal effects of the interaction between year and neural network complexity for the log RMSE ratio produced using the marginaleffects R package (Arel-Bundock et al.,, 2024). Predicted values and 95% confidence bands were calculated using the empirical distribution of all other covariates, providing a marginalised view over the observed data. This highlights how the performance of GAMs relative to neural networks changes over time across different complexity levels (where, again, positive values indicate that neural networks perform better). While performance for small and large networks remains relatively stable, medium complexity neural networks show an improvement over time. In contrast, datasets using neural networks of unknown complexity in earlier years lie just above the zero line, and this trend declines slightly over time. For all complexities in 2024, the confidence bands include zero, indicating no clear performance advantage of any neural network complexity over GAMs at present. Rug marks along the x-axis indicate the years for which there are observations for each complexity level, highlighting where data are available and model estimates are most reliable.
@@
No_Predictorsmedium $\phantom{-}0.26$ (0.23) -0.08 (0.24)  0.34 (0.43)
No_Predictorslarge  0.35 (0.34)  2.19 (0.70)  0.50 (0.82)
No_Predictorsunknown  0.11 (0.38)  0.89 (1.04)  0.25 (0.78)
@@
Year:Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.02 (0.06)
Year:Sample_Sizelarge -0.06 (0.04) -0.29 (0.05) -0.07 (0.12)
Year:Sample_Sizeunknown -0.10 $(0.03)^{*}$ -0.21 (0.09)  0.00 (0.08)
Year:No_Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year:No_Predictorslarge -0.03 (0.04) -0.25 (0.09) -0.06 (0.08)
Year:No_Predictorsunknown -0.01 (0.04) -0.15 (0.13)  0.00 (0.08)
Year:Complexitymedium  0.11 (0.04)  0.04 (0.07)
Year:Complexitylarge  0.10 (0.05)  0.01 (0.08)
Year:Complexityunknown  0.01 (0.03) -0.06 (0.13) -0.09 (0.05)
@@
From the model output, the datasets of unknown sample size have a positive main effect on the log RMSE ratio ($\hat{\beta}$ = 1.25, SE = 0.35) and a negative interaction with year ($\hat{\beta}$ = 0.10-0.10, SE = 0.03). Thus, at the 2011 baseline, neural networks outperform GAMs for datasets with unknown sample size relative to the reference category (small). However, the advantage significantly reduces over time; in any case, there is little practical implication here, as it is unclear how the datasets of unknown size differ from those where the sample size is reported.
The ICC is high (0.77), indicating substantial between-paper variability and supporting the appropriateness of a mixed-effects model. The addition of covariates to the model does not explain the variation across papers.
7.1.3 AUC metric analysis
The AUC analysis included 111 observations from 37 papers. No outliers were identified, and all data points were included in the analysis. For neural network complexity, the large category was not represented in the data and hence omitted from the model. Small and medium categories were merged due to the minimal observations in each, and hence small/medium was used as the reference level.

### v2

@@ 6.2 Dataset-level characteristics
MLP neural networks were the most common type, accounting for 89% of datasets. All other neural networks were variations of this, and any one type accounted for less than $5\%$ of the datasets. Neural network complexity was calculated as the total number of parameters, excluding those in the input layer. This exclusion was necessary because input nodes correspond to the number of predictors, which is already captured by the fact that we have collected the number of predictor variables as an attribute. As mentioned, neural network complexity is among the data items frequently unreported. Only 35% of datasets provided sufficient detail to compute the full parameter count. However, 64% reported the number of hidden layers, which includes instances where layers were discussed but not the number of nodes in each needed to look at full complexity. Depth of a neural network via the number of hidden layers is, in its own right, a feature of interest. Among the datasets that did report the number of hidden layers, 70% used a single hidden layer, 22% had two hidden layers, and 7% used between three and five layers. This distribution indicates that the majority of neural networks used in comparisons with GAMs were shallow architectures, typically with one or two hidden layers.
@@
Looking at author-reported outcomes gives us an insight into RQ2, which investigates which model is better overall. Neural networks were declared better than GAMs in 27% of datasets. This is more than the 19% of datasets in which GAMs were reported as better than neural networks. In 9% of datasets, both model types were found equivalent. However, author-reported outcomes provide an incomplete and often inconsistent comparison between GAMs and neural networks. Differences can arise in how outcomes are classified from author to author. Additionally, 46% of author outcomes are undefined and hence recorded as “unclear”, often due to a focus on an alternative model in that application. A way to address these limitations is to directly analyse the performance metric values reported for both GAMs and neural networks in each instance. This allows for a more objective comparison and quantifies the magnitude of performance differences, which is not possible when looking at the qualitative author-reported outcomes alone; it also provides a firmer, more quantitative answer to RQ2.
@@ 7 Performance metric analysis
To quantitatively synthesize the comparative performance of GAMs and neural networks across the reviewed literature, we create a statistical model for the reported performance metric values directly. This analysis examines how the relative performance, measured by the most frequently used metrics, is associated with key study characteristics such as sample size, domain area, and model complexity. This will address RQ2 and RQ3. RQ3 examines whether or not available data or modelling characteristics can explain observed differences in model performance. To answer these, we model the reported performance metrics from the 430 datasets in terms of the available characteristics. Given the hierarchical structure of the data we have collected, where datasets are nested within papers, a mixed-effects modelling approach was employed to account for intra-paper dependence.
@@ 7.1 Methods
Linear mixed-effects models were fitted using the lme4 package (lme4_cite) in R programming language (R_cite). We included a paper-level random intercept effect to account for correlated results within a given paper, along with the following fixed effects: publication year, domain area, sample size, number of predictors, and neural network complexity. Pairwise interactions with year of publication were also included to account for trends over time.
7.2 Response and predictor variables
As shown in Table 8, RMSE, $R^{2}$, and AUC are, by far, the most commonly-used performance metrics, accounting for 53%, 33%, and 26% of datasets respectively and collectively covering 89% of all datasets. RMSE and $R^{2}$ are common performance metrics used in regression, accounting for 76% and 48% of regression datasets respectively, and 91% collectively. AUC, a standard metric for classification models, was used in 82% of the classification datasets. These three metrics were selected for our performance metric analysis due to their prevalence and large coverage of datasets.
@@ 7.2 Response and predictor variables
where positive values indicate superior neural network performance for that metric, negative values signify GAM performance was better, and zero indicates model equivalence. For AUC and $R^{2}$, they are first converted to losses, in which 0 indicates optimal performance, by subtracting them from 1. This allows for all three to represent loss ratios. The ratio of performance metrics provides a relative performance measure that is invariant to scale. This is necessary here, as we are comparing results across many different datasets with substantial variation in absolute performance; different datasets lie on different scales, with differing levels of achievable performance given the inherent noise/uncertainty in the problem at hand. Moreover, we have taken the logarithm to place the response on the real line (so that it is suitable for linear modelling) and to address some skewness. Figure 5 displays the log performance ratios for datasets using RMSE, AUC, and $R^{2}$; points are coloured based on the performance metric used. The horizontal line at zero indicates where GAMs and neural networks are equivalent; points above the line correspond to datasets where neural networks outperform GAMs, while points below are where GAMs achieve superior performance. Overall, we can see that there is a high level of variability in results, with no firm evidence in favour of neural networks or additive models (as the points are generally centred around zero), and no strong trends over time. Thus, while neither model class appears to be systematically superior across the studies (addressing RQ2), this high-level overview of results does not account for study characteristics (other than the year) or the random effects of papers. These additional aspects are the focus of RQ3, and will be considered throughout the remainder of this paper.
@@
The following study characteristics (i.e., paper and dataset characteristics) were considered in the meta-analysis as fixed effects: year of publication (Year), domain area (Area), sample size (Sample_Size), number of predictors (Predictors), and neural network complexity (Complexity). Paper ID (Paper_ID) is used as a random effect to account for the dependence between datasets coming from the same papers. To retain as much information as possible, the numerical variables with missing values (described in Section 5.1) were converted to categorical variables. Categories were formed by splitting the numerical variables by their observed tertiles, along with a further “Unknown” category for those with missing values, which resulted in four levels. This approach enabled the inclusion of incomplete records and allowed assessment of whether the presence of missingness itself was associated with performance outcomes. A detailed overview of all variables included in our meta analysis model for RQ3 is provided in Table 9.
Table 9: Summary of response variable, predictor variables, transformations, and associated boundary conditions used in the meta analysis model.
Component & Type Description & boundary conditions
Response variables
Log performance ratio                 Continuous (real number) Defined as $\log(\mathrm{RMSE}_{\mathrm{GAM}}/\mathrm{RMSE}_{\mathrm{NN}})$ $\log\{(1-R^{2}_{\mathrm{GAM}})/(1-R^{2}_{\mathrm{NN}})\}$ $\log\{(1-\mathrm{AUC}_{\mathrm{GAM}})/(1-\mathrm{AUC}_{\mathrm{NN}})\}$ Note that $R^{2}$ and AUC were first converted to losses ($1-R^{2}$ and $1-\mathrm{AUC}$) to align with RMSE, and the log transformation places the responses on the real line such that positive values imply that NNs are better and negative values imply that GAMs are better.
Predictor variables (fixed effects)
Year                 Numerical (integer) Year of study (1998 to 2024) centred by subtracting the middle year, 2011 (leading to the range 13-13 to $13$).
Area                 Categorical (4 levels) Domain area classified from Scopus categories as: Environmental and Agricultural Sciences; Engineering, Mathematics and Computer Science; Health and Biomedical Sciences; and Other.
@@
Sample_Size                 Categorical (4 levels) Number of observations in dataset: Small $\leq 350$, Medium (35035028002800), Large $\geq 2800$, and Unknown.
@@
Predictors                 Categorical (4 levels) Number of predictors in dataset: Small $\leq 5$, Medium (551010), Large $\geq 10$, and Unknown.
@@
Complexity                 Categorical (4 levels) Number of neural network parameters (excluding input layer): Small $\leq 6$, Medium (66150150), Large $\geq 150$, and Unknown.
Random effect
Paper_ID                 Categorical (143 levels) The ID of the paper, which accounts for correlation of multiple datasets originating from that paper.
Additional modelling details
Pairwise interactions Interactions between Year and the other predictors were fitted to investigate whether effects changed over time.
Reference levels For Year: 2011 (centred to be 0). For categorical variables Sample_Size, Predictors, and Complexity: “Small”.
We consider the linear mixed–effects model with random intercepts:
$Y_{ij}=\beta_{0}+\alpha_{j}+X_{ij}\beta+\varepsilon_{ij},\qquad\alpha_{j}\sim N(0,\sigma_{\alpha}^{2}),\quad\varepsilon_{ij}\sim N(0,\sigma^{2}),$ (7.7)
where $i$ indicates datasets within paper $j$. Here $\beta_{0}$ and $\beta$ are fixed effects, while $\alpha_{j}$ represents the paper‐specific random intercept capturing paper-level heterogeneity. The residual term $\varepsilon_{ij}$ represents dataset-level noise, assumed independent given the fixed and random effects. With the notation of (7.7) in place, note that the intraclass correlation, described Section 7.1, is given by $\mathrm{ICC}=\sigma_{\alpha}^{2}/(\sigma_{\alpha}^{2}+\sigma^{2})$. Figure 6 provides a graphical representation of the random-intercept model in (7.7). The outer box corresponds to a paper $j$, and the inner box represents a dataset $i$ within that paper. The random intercept $\alpha_{j}$ is a paper-level effect that influences all outcomes $Y_{ij}$ in that paper. The arrow from $\alpha_{j}$ to $Y_{ij}$ is the shared paper-specific shift in the mean response. The relationship between $X_{ij}$, $Y_{ij}$, and any unobserved latent variables $H_{ij}$ is shown inside the dataset box.
Figure 6: Linear mixed-effects model diagram where $Y_{ij}$ is the response for the RMSE model case. For the $R^{2}$ model the response is $\log\{(1-R^{2}_{\mathrm{GAM}})/(1-R^{2}_{\mathrm{NN}})\}$ and AUC model response is $\log\{(1-\mathrm{AUC}_{\mathrm{GAM}})/(1-\mathrm{AUC}_{\mathrm{NN}})\}$. The random intercept is denoted by $\alpha_{j}$, $X_{ij}$ are the set of fixed-effects, and $H_{ij}$ are any potential latent variables.
7.2.1 Unrecorded variables
Several additional potentially important dataset-level variables were not consistently reported within the reviewed literature, including data preprocessing steps, optimisation algorithms and settings, software environments, neural network hyperparameter tuning procedures, activation functions, and GAM details (e.g., spline type, degrees of freedom, knot placement, and smoothness penalty). These variables can influence model performance and therefore represent potential confounding variables in comparative analyses. However, because most papers did not report these in sufficient detail, they could not be incorporated as variables in this analysis. The absence of such information likely introduces additional heterogeneity that cannot be modelled directly, but may be captured to some extent by the paper random effects included in our meta analysis model.
This limitation reflects a broader reporting issue common within machine learning research. There is an absence of standardised reporting practices for model development and evaluation. As highlighted in mitchell2019model, there is a growing recognition that inadequate documentation obscures methodological choices and hinders reproducibility of results. Our systematic review strongly reinforces this previous research as we find that unreported information is prevelant across a large number of papers, from different fields, published over a 25-year period. Thus, there is a need for more rigorous reporting protocols even in simpler tabular dataset modelling.
Due to the unreported variables, the effects in the following sections are interpreted as associations rather than as causal mechanisms. Moreover, missing information is a common feature of systematic reviews in general as discussed previously in Section 5.1. In any case, it is of interest to explore the associations between model performance and key variables (e.g., year of study, discipline, data dimensions) per RQ3 within this systematic review, which is the first review of papers comparing neural networks and GAMs.
7.3 RMSE metric analysis
@@
The RMSE analysis included 229 dataset observations from 80 papers. After removing two outliers, 227 dataset observations, from 78 papers remained. Including covariates in the analysis allows us to explain more of the variation in RMSE ratios across datasets, and to investigate how aspects of the datasets and models are associated with the relative performance of neural networks and GAMs. The RMSE model estimates and standard errors (SE) are presented in Table 10. (Models for AUC and $R^{2}$ are also shown, but discussed in Section 7.4 and Section 7.5 respectively.) The model included all fixed-effects predictor variables and their pairwise interactions with time (Year) as follows:
@@
where $(1|\texttt{Paper\_ID})$ represents a random intercept capturing variation among datasets. Note that modelling the ratio in this manner is a special case of a slightly more general model discussed in A.
@@
Predictorsmedium $\phantom{-}0.26$ (0.23) -0.08 (0.24) 0.34 (0.43)
Predictorslarge 0.35 (0.34) 2.19 (0.70) 0.50 (0.82)
Predictorsunknown 0.11 (0.38) 0.89 (1.04) 0.25 (0.78)
@@
Year × Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.02 (0.06)
Year × Sample_Sizelarge -0.06 (0.04) -0.29 (0.05) -0.07 (0.12)
Year × Sample_Sizeunknown -0.10 $(0.03)^{*}$ -0.21 (0.09) 0.00 (0.08)
Year × Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year × Predictorslarge -0.03 (0.04) -0.25 (0.09) -0.06 (0.08)
Year × Predictorsunknown -0.01 (0.04) -0.15 (0.13) 0.00 (0.08)
Year × Complexitymedium 0.11 (0.04) 0.04 (0.07)
Year × Complexitylarge 0.10 (0.05) 0.01 (0.08)
Year × Complexityunknown 0.01 (0.03) -0.06 (0.13) -0.09 (0.05)
@@
The intercept for this model is positive, however, not statistically significantly different from zero. This indicates that, at the reference levels for covariates, there is no evidence of a distinction between the performance of GAMs and neural networks (as already noted in Figure 5). Neural network complexity and its time interaction are statistically significant. Specifically, medium and large complexity effects are both negative and significant, indicating that, in 2011, more complex neural networks performed worse than GAMs compared to smaller neural networks. The complexity-year interactions for medium and large neural networks are positive and significant; this indicates that larger neural networks are associated with improved performance over time relative to small neural networks.
@@
Figure 7 displays the marginal effects of the interaction between year and neural network complexity for the log RMSE ratio produced using the marginaleffects R package (mixedeffectsR). Predicted values and 95% confidence bands were calculated using the empirical distribution of all other covariates, providing a marginalised view over the observed data. This highlights how the performance of GAMs relative to neural networks changes over time across different complexity levels (where, again, positive values indicate that neural networks perform better). The main finding is that the medium complexity neural networks are associated with worse performance relative to GAMs, but the effect is non-significant in the most recent five years. Indeed, for all complexity levels in 2024, the confidence bands include zero, indicating that neural network complexity is not associated with performance relative to GAMs at present (albeit other unreported neural network or GAM details could be). Rug marks along the x-axis indicate the years for which there are observations for each complexity level, highlighting where data are available and model estimates are most reliable.
We also see from Table 10 that datasets of unknown sample size have a positive main effect ($\hat{\beta}$ = 1.25, SE = 0.35) and a negative interaction with year ($\hat{\beta}$ = 0.10-0.10, SE = 0.03), both statistically significant. Thus, at the 2011 baseline, neural networks appear to outperform GAMs in datasets with unknown sample size relative to the reference category (small), with this effect reducing over time. However, there is little practical implication here, as it is unclear how the datasets of unknown size differ from those where the sample size is reported.
The ICC is high (0.77), indicating substantial between-paper variability and supporting the appropriateness of a mixed-effects model in our analysis. It is possible that the large variability between papers could be explained by unreported features within papers that may influence model performance, e.g., data preprocessing and optimisation settings (see Section 7.2.1).
7.4 AUC metric analysis
The AUC analysis included 111 observations from 37 papers. No outliers were identified, and all data points were included in the analysis. For neural network complexity, the “Large” category was not represented in the data and hence omitted from the model. The “Small” and “Medium” categories were merged due to the minimal observations in each, and hence “Small/Medium” was used as the reference level.